OpenStudy (anonymous):

] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (anonymous):

substitution: i feel like it 4x + 3y = -1 3x + y = 3 ----->y = 3-3x 4x + 3y = -1 4x + 3(3-3x) = -1 4x + 9-9x = -1 -5x + 9 = -1 -5x = -10 x = 2 y = 3-3x y = 3-3(2) y = 3-6 y = -3

OpenStudy (anonymous):

THANK YOU!